OpenStudy (anonymous):

what is the simplest form of a^9b^5/ a^3b^15

jhonyy9 (jhonyy9):

a6b(-10)

jimthompson5910 (jim_thompson5910):

Subtract corresponding exponents to get a^(9-3)b^(5-15) = a^6b^(-10) = (a^6)/(b^10) So the answer is \[\Large \frac{a^{6}}{b^{10}}\]

OpenStudy (anonymous):

thanks!